OpenStudy (anonymous):

At 95% humidity, what percentage of the air around us is water vapour? Can you drown?

OpenStudy (king):

at any humidity i dont think we can drown bcuz at an high humidity we will sweat a lot and thats all and our sweat is mainly salt water so we cant drown unless we push ourselves down and suicide!

OpenStudy (anonymous):

the humidity means the percentage of water mass in the total weight of steam,therefore the mass percentage of air is 5%. we die from Suffocation

OpenStudy (jamesj):

There are two types of Humidity: Absolute Humidity and Relative Humidity. Absolute Humidity is the weight of water vapor in a cubic meter of air. The higher the temperature, the higher the ability of air to hold water. For example, at 20 C, the maximum Absolute Humidity is 17.3 g/m^3. At 50 C, the maximum Absolute Humidity is 83.0 g/m^3 Relativity Humidity is the % of water vapor in the air relative to the maximum possible Absolute Humidity. For example, if the air temperature is 20 C and the Absolute Humidity is 10.0 g/m^3, then the Relative Humidity is \[ Relative \ Humidity=\frac{10.0}{17.3}=0.58=58\% \] When we use the word Humidity is ordinary speech, we're talking about Relative Humidity. So ... if the humidity is 95% or even 100%, that doesn't mean 95% or 100% of the air around us is water vapor; just the amount of water in the air is 95% or 100% of the maximum it can hold. Even at 50 C and 100% Relative Humidity, there is only 83.0 grams per cubic meter. Now 1 cubic meter = (100)^3 cubic cm = 1,000,000 cubic cm If you collected all 83 g of the water together, as 1 gm of water = 1 cm then 83.0 grams = 83 cubic cm Hence in 1 cubic cm, 83 g of water is 83/1,000,000 = 0.0083% much less than 1% So even at 100% Humidity at 50 C we wouldn't drown. We'd be very uncomfortable, from the heat content of the air and water, but we could still breath.